Suppose a two-period-lived household facing a financial market in which it can borrow or lend at an interest rate equal to rThe correct answers are in bold. I need help seeing the derivation of these answers from the given problem. Thank you! Income is exogenous and given in the beginning of problem (denoted by Y1 & Y2)
Answer #1

C Page No : at Eqm, 80 from MRS1,2= (1+0) C2 = {B (HO)] / *a B.C. 1.15a + Cq = (1+8) Yi+Y 2 = 4.15 1159 + 9• (-95 (1•15)3 =d) at Eqm c = [puto) ] Vi 4G (a) = B lito) (to)= (2) l of 10- (Text) of -1 Ang (d) . Yt